OpenStudy (anonymous):

Could somebody please help me show my work? I have the right answer the question, I just need to show how I got that answer.. 1. Which values of a and b are a solution to the inequality I 5 - 2a I - B < = 4? Answer: a = 3, b = -1

OpenStudy (anonymous):

for this problem there are an infinite number of solutions the only thing you can do is check your possible answers

OpenStudy (anonymous):

\[|5-2a|-b\leq 4\] \[|5-2a|\leq 4-b\] \[b-4\leq 5-2a\leq 4-b\] is a start if you try to solve it using algebra but you have two unknowns

OpenStudy (anonymous):

i am assuming you were given some choices and asked "which ones work" or something like that

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

then all you can do is check them plug in the numbers and see which ones work

OpenStudy (anonymous):

I know the answer, I just need to show my work, im bad at showing my work /: